### v1

@@ Results
To further probe the robustness of the main results, we estimate a set of distributional regressions. Rather than focusing solely on the continuous homicide rate, this approach assesses whether environmental enforcement affects the likelihood that a municipality experiences any violence or presents a pattern of persistently high violence over time. We define “any violence” as a binary indicator equal to one if the homicide rate is greater than zero, and endemic violence as a binary indicator equal to one if the homicide rate exceeds 10 homicides per 100,000 inhabitants in a given year. The latter threshold draws from public health literature, where homicide rates above 10 are often considered a marker of endemic violence by institutions such as the World Health Organization (WHO).
@@ Conclusion
The implications of these findings are significant for policymakers addressing governance challenges in environmentally sensitive areas. Traditionally, policy discussions around conservation and development have been framed as involving trade-offs, suggesting that strengthened environmental regulation could constrain economic opportunities and exacerbate social tensions. However, our results challenge this conventional wisdom by demonstrating that command-and-control policies designed to prevent illegal deforestation can simultaneously mitigate violent crime. Consequently, policymakers should integrate environmental enforcement measures into broader public safety and governance strategies, especially in regions characterized by institutional fragility and limited accessibility.
@@
Appendix A Appendix Crime, Environment, and Governance
@@
These patterns are also reflected in the data used in this study. Figure 1 shows the evolution of homicide rates in municipalities located inside and outside the Amazon biome between 2000 and 2016.666Section 4 details on the construction of the homicide rate variable. At the beginning of the period, municipalities outside the biome had higher rates of lethal violence. However, trajectories diverged in the early 2000s, as homicide rates in Amazon municipalities began to rise more rapidly, while rates in non-Amazon areas remained relatively stable. By 2016, this divergence had become substantial: average homicide rates in the Amazon far exceeded those in the rest of the country. This trend reinforces the idea that institutional fragility and the expansion of illegal markets have created a distinct environment marked by growing and persistent violence in the region.
Between 2006 and 2016 — the core period of this study — the average homicide rate in Amazon municipalities rose from 33.1 to 52.1 per 100,000 inhabitants, a 57.3% increase. In contrast, non-Amazon municipalities experienced only a modest rise of 8.0%, from 36.8 to 39.8. To ensure that the striking pattern of rising and comparatively higher homicide rates in the Amazon is not merely the result of aggregation masking regional heterogeneity, we also disaggregated the data by Brazil’s official geographic regions. This additional breakdown further supports the main result. As shown in Table A.1, the North region recorded the sharpest increase in homicide rates over the period, rising from 33.1 in 2006 to 53.5 in 2016, a 61.5% growth.777 The slight difference between this two growth rates reflects the fact that not all municipalities in the North region are part of the Amazon biome, and not all Amazon biome municipalities belong to the North region. Together, the figure and the table provide empirical confirmation of the institutional dynamics discussed above, highlighting the concentration of violence in the Amazon region.
@@
Most of the deforestation observed in the Brazilian Amazon over the past two decades has been illegal (Assunção et al., 2023). Under the Brazilian Forest Code,888 The Brazilian Forest Code is a set of federal laws (primarily Law No. 12.651/2012) regulating the use and protection of native vegetation on private rural properties in Brazil. legal deforestation is permitted only within defined limits and under strict regulation — for example, inside rural private properties and with formal authorization. However, even in cases where deforestation could be legal in principle, property-level assessments reveal widespread noncompliance with environmental laws (Godar et al., 2012). Moreover, a significant portion of forest clearing occurs in areas where all deforestation is prohibited, such as protected lands, indigenous territories, and state-owned undesignated forests (Moutinho and Azevedo-Ramos, 2023). In these cases, deforestation frequently serves as a strategy to assert possession, particularly in undesignated lands, where demonstrating productive use can strengthen claims for future land titles (Aldrich et al., 2012; Brown et al., 2016).
@@
At the core of the PPCDAm is the satellite-based Real-Time Detection of Deforestation System (DETER), developed and operated by INPE. The DETER system processes satellite imagery to rapidly identify signs of forest cover change, issuing georeferenced alerts that guide inspection teams in the field. Between 2004 and 2017, DETER issued more than 70,000 alerts, corresponding to approximately 88,000 km² of forest disturbance (INPE, 2018). DETER initially relied on MODIS sensor data from NASA’s Terra satellite, with a spatial resolution of 250 meters and daily coverage. This configuration allowed for frequent updates, though it limited detection to forest changes larger than 25 hectares. Analysts at INPE visually interpret imagery using spectral mixture models and predefined classification rules, identifying various categories of disturbance: clear-cut deforestation (with or without remaining vegetation), forest degradation (including fire scars), and logging activity.999 Detailed methodological documentation is provided in INPE (2022a).
@@
Evidence shows that DETER has been effective in curbing deforestation. Assunção et al. (2023) document a significant reduction in forest loss inside the Amazon biome following the introduction of real-time satellite monitoring.101010 Although DETER monitoring formally covers the full Legal Amazon, 97% of the area deforested since its creation has been concentrated within the Amazon biome (INPE, 2022b). Beyond environmental outcomes, enforcement efforts may also affect broader dimensions of state presence and local behavior. As noted by Dechezleprêtre and Sato (2017), environmental policies often generate spillover effects on institutional quality, economic incentives, and conflict dynamics. In this sense, DETER serves not only as a monitoring tool but also as a mechanism through which the state reasserts its authority in contested territories.
@@ Appendix B Data Construction
Despite these limitations, the Ministry of Health estimates indicate that SIM captures over 96% of deaths nationwide, and more then 90% in the North region (Brazilian Ministry of Health, 2011). This high coverage makes SIM reliable for population-level analyses (Cunha et al., 2017). Mathers et al. (2005) also highlight the quality of SIM data not only for the count of deaths, but also for the consistent use of the International Statistical Classification of Diseases and Related Health Problems, 10th Revision (ICD-10).111111 The ICD-10 is maintained by the World Health Organization and provides a standardized international framework for coding mortality and morbidity data (World Health Organization, 2019).
@@ B.4.1 Commodity Price Index
Appendix C Controlling for Conservation Policies

### v2

@@ Results
To further probe the robustness of the main results, we estimate a set of distributional regressions. Rather than focusing solely on the continuous homicide rate, this approach assesses whether environmental enforcement affects the likelihood that a municipality experiences any violence or presents a pattern of persistently high violence over time. We define “any violence” as a binary indicator equal to one if the homicide rate is greater than zero, and endemic violence as a binary indicator equal to one if the homicide rate exceeds 10 homicides per 100,000 inhabitants in a given year. The latter threshold draws from the public health literature, where homicide rates above 10 are often considered a marker of endemic violence by institutions such as the World Health Organization (WHO).
@@ Conclusion
The implications of these findings are significant for policymakers addressing governance challenges in environmentally sensitive areas. Traditionally, policy discussions around conservation and development have been framed as involving trade-offs, suggesting that strengthened environmental regulation could constraint economic opportunities and exacerbate social tensions. However, our results challenge this conventional wisdom by demonstrating that command-and-control policies designed to prevent illegal deforestation can simultaneously mitigate violent crime. Consequently, policymakers should integrate environmental enforcement measures into broader public safety and governance strategies, especially in regions characterized by institutional fragility and limited accessibility.
@@
Appendix A Detailed Institutional Background: Crime, Environment, and Governance
@@
These patterns are also reflected in the data used in this study. Figure 1 shows the evolution of homicide rates in municipalities located inside and outside the Amazon biome between 2000 and 2016.111111Section 4 details on the construction of the homicide rate variable. At the beginning of the period, municipalities outside the biome had higher rates of lethal violence. However, trajectories diverged in the early 2000s, as homicide rates in Amazon municipalities began to rise more rapidly, while rates in non-Amazon areas remained relatively stable. By 2016, this divergence had become substantial: average homicide rates in the Amazon far exceeded those in the rest of the country. This trend reinforces the idea that institutional fragility and the expansion of illegal markets have created a distinct environment marked by growing and persistent violence in the region.
Between 2006 and 2016 — the core period of this study — the average homicide rate in Amazon municipalities rose from 33.1 to 52.1 per 100,000 inhabitants, a 57.3% increase. In contrast, non-Amazon municipalities experienced only a modest rise of 8.0%, from 36.8 to 39.8. To ensure that the striking pattern of rising and comparatively higher homicide rates in the Amazon is not merely the result of aggregation masking regional heterogeneity, we also disaggregated the data by Brazil’s official geographic regions. This additional breakdown further supports the main result. As shown in Table A.1, the North region recorded the sharpest increase in homicide rates over the period, rising from 33.1 in 2006 to 53.5 in 2016, a 61.5% growth. [^12] The slight difference between this two growth rates reflects the fact that not all municipalities in the North region are part of the Amazon biome, and not all Amazon biome municipalities belong to the North region. Together, the figure and the table provide empirical confirmation of the institutional dynamics discussed above, highlighting the concentration of violence in the Amazon region.
@@
Most of the deforestation observed in the Brazilian Amazon over the past two decades has been illegal (Assunção et al., 2023). Under the Brazilian Forest Code, [^13] The Brazilian Forest Code is a set of federal laws (primarily Law No. 12.651/2012) regulating the use and protection of native vegetation on private rural properties in Brazil. legal deforestation is permitted only within defined limits and under strict regulation — for example, inside rural private properties and with formal authorization. However, even in cases where deforestation could be legal in principle, property-level assessments reveal widespread noncompliance with environmental laws (Godar et al., 2012). Moreover, a significant portion of forest clearing occurs in areas where all deforestation is prohibited, such as protected lands, indigenous territories, and state-owned undesignated forests (Moutinho and Azevedo-Ramos, 2023). In these cases, deforestation frequently serves as a strategy to assert possession, particularly in undesignated lands, where demonstrating productive use can strengthen claims for future land titles (Aldrich et al., 2012; Brown et al., 2016).
@@
At the core of the PPCDAm is the satellite-based Real-Time Detection of Deforestation System (DETER), developed and operated by INPE. The DETER system processes satellite imagery to rapidly identify signs of forest cover change, issuing georeferenced alerts that guide inspection teams in the field. Between 2004 and 2017, DETER issued more than 70,000 alerts, corresponding to approximately 88,000 km² of forest disturbance (INPE, 2018). DETER initially relied on MODIS sensor data from NASA’s Terra satellite, with a spatial resolution of 250 meters and daily coverage. This configuration allowed for frequent updates, though it limited detection to forest changes larger than 25 hectares. Analysts at INPE visually interpret imagery using spectral mixture models and predefined classification rules, identifying various categories of disturbance: clear-cut deforestation (with or without remaining vegetation), forest degradation (including fire scars), and logging activity. [^14] Detailed methodological documentation is provided in INPE (2022).
@@
Evidence shows that DETER has been effective in curbing deforestation. Assunção et al. (2023) document a significant reduction in forest loss inside the Amazon biome following the introduction of real-time satellite monitoring. Beyond environmental outcomes, enforcement efforts may also affect broader dimensions of state presence and local behavior. As noted by Dechezleprêtre and Sato (2017), environmental policies often generate spillover effects on institutional quality, economic incentives, and conflict dynamics. In this sense, DETER serves not only as a monitoring tool but also as a mechanism through which the state reasserts its authority in contested territories.
@@ Appendix B Data Construction
Despite these limitations, the Ministry of Health estimates indicate that SIM captures over 96% of deaths nationwide, and more then 90% in the North region (Brazilian Ministry of Health, 2011). This high coverage makes SIM reliable for population-level analyses (Cunha et al., 2017). Mathers et al. (2005) also highlight the quality of SIM data not only for the count of deaths, but also for the consistent use of the International Statistical Classification of Diseases and Related Health Problems, 10th Revision (ICD-10). [^15] The ICD-10 is maintained by the World Health Organization and provides a standardized international framework for coding mortality and morbidity data (World Health Organization, 2019).
@@ B.4.1 Commodity Price Index
Appendix C Heterogeneity by Death Category
In this appendix, we explore the robustness of our results to alternative and more disaggregated homicide classifications. Specifically, we consider (i) assaults committed with firearms (ICD-10 X93–X95), (ii) assaults committed with sharp or blunt objects (ICD-10 X99–Y00), (iii) deaths resulting from legal intervention (ICD-10 Y35), and (iv) deaths resulting from intentional self-harm (ICD-10 X60–X84). These categories are motivated by the literature and by concerns raised in previous discussions: firearm-related deaths, assaults with physical objects and intentional self-harm capture different types of violence, while legal intervention provides a measure of fatalities caused by state agents in contexts where lethal force is officially sanctioned. Although enforcement in the Amazon is not primarily conducted by conventional police forces but by environmental agencies such as IBAMA, these distinctions allow us to test whether changes in enforcement affect specific causes of violent mortality.
Table C.1 reports the 2SLS estimates for these alternative outcomes. The results indicate no statistically significant association between environmental enforcement and firearm-related homicides (Column (1)). By contrast, the coefficient for assaults with sharp or blunt objects (Column (2)) is negative and statistically significant at the 5% level: one additional fine is associated with a reduction of 0.33 homicides per 100,000 inhabitants, corresponding to a 3.7% decrease relative to the average homicide rate in this category. This effect is sizeable, implying that moving from the 25th to the 75th percentile in fines (roughly nine additional fines) is associated with a reduction of about one-third of the average homicide rate due to assaults with sharp or blunt objects. Additionally, the coefficient for deaths due to legal intervention (Column (3)) is small and statistically insignificant, suggesting that increased environmental enforcement is not associated with higher rates of fatalities in confrontations with state agents. This result is consistent with the institutional context in which IBAMA, rather than police forces, carries out most enforcement activities, and confrontations rarely escalate into lethal violence. Lastly, the coefficient for deaths due to intentional self-harm (Column (4)) is small and statistically insignificant, suggesting that increased environmental enforcement is not associated with higher rates of suicides.
Table C.1: 2SLS — Estimates under Alternative Homicide Classifications
Homicide Rate
(1) (2) (3) (4)
Lagged Enforcement -0.132 -0.330∗∗ 0.002 -0.048
(0.194) (0.165) (0.008) (0.078)
Average homicide rate across municipalities 11.65 8.89 0.04 1.95
FE: Municipality & Year
All controls
Observations 5,210 5,210 5,210 5,210
Municipalities 521 521 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) focuses on assault by firearm (ICD-10 X93–X95), Column (2) focuses on assault by sharp or blunt object (ICD-10 X99–Y00), Column (3) focuses on legal intervention (ICD-10 Y35), and Column (4) focuses on intentional self-harm (ICD-10 X60–X84). “Homicide Rate” is the number of homicides per 100,000 inhabitants. “Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006–2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
Appendix D Heterogeneity by Sex, Race and Age Group
A natural question is whether the estimated effects of environmental enforcement on homicides may vary across demographic groups. For example, BFPS and IPEA (2025) show national evidence that lethal violence is disproportionately concentrated among young men and African Brazilians.
In line with IBGE (Brazilian Census Bureau) classification, we group individuals into three broad racial categories: White and Yellow, African Brazilian (Black and Pardo), and Indigenous. The African Brazilian group combines individuals classified as “Preto” (Black) and Pardo. While “Preto” refers to individuals self-identified as Black, Pardo is a census category used in Brazil to capture mixed-race individuals with diverse backgrounds, reflecting the country’s history of racial mixture (IBGE, 2013). We adopt the term African Brazilian to designate this combined group throughout the analysis.
In this appendix, we analyze differences by sex (Men v. Women), race (White and Yellow v. African Brazilian v. Indigenous) and age groups (0–15 years v. 15–25 years v. 25–35 years v. 36+ years). To this end, we divide this appendix into three sections: (i) sex, (ii) race and (iii) age groups. In each section, we first present descriptive statistics on the distribution of homicides and then estimate the same regression specification as in our main analysis.
Sex
Table D.1 reports the annual distribution of homicide victims by sex between 2006 and 2016. The descriptive statistics reveal that the vast majority of homicide victims are men, consistently representing around 87–88% of all cases throughout the period, while women account for roughly 12% and cases of unknown sex remain negligible. This stability over time is in line with the broader literature documenting that lethal violence in Brazil is overwhelmingly concentrated among men (BFPS and IPEA, 2025).
Table D.1: Annual Totals and Shares of Deaths by Sex
Year Men Women Unknown Total Share of Share of Share of
men (%) women (%) unknown (%)
(1) (2) (3) (4) (5) (6) (7) (8)
2006 59,068 7,965 90 67,123 88.0 11.9 0.1
2007 61,235 8,412 92 69,739 87.8 12.1 0.1
2008 63,926 8,900 92 72,918 87.7 12.2 0.1
2009 64,505 9,087 95 73,687 87.5 12.3 0.1
2010 62,160 9,155 78 71,393 87.1 12.8 0.1
2011 64,456 9,436 101 73,993 87.1 12.8 0.1
2012 67,605 9,474 119 77,198 87.6 12.3 0.2
2013 69,977 9,404 142 79,523 88.0 11.8 0.2
2014 68,546 9,481 125 78,152 87.7 12.1 0.2
2015 71,921 9,546 137 81,604 88.1 11.7 0.2
2016 75,546 9,690 151 85,387 88.5 11.3 0.2
Notes: Columns (2)–(4) report annual totals of deaths for each sex category. Column (5) reports the total number of deaths across these three categories. Columns (6)–(8) present the share of each sex category relative to the total for the corresponding year, expressed in percent. Source: SIM-DataSUS
Table D.2 presents the 2SLS estimates by sex. The results indicate that environmental enforcement significantly reduces homicides among men: the coefficient is –0.75 and statistically significant at the 5% level. Given an average male homicide rate of 25.05 per 100,000 inhabitants, this estimate implies that the issuance of one additional fine is associated with a reduction of about 3.0% relative to the mean male homicide rate. By contrast, the effect for women is close to zero (0.005) and not statistically significant, consistent with the expectation that environmental enforcement primarily affects types of violence linked to illegal economic activities, which disproportionately involve men (Soares et al., 2021; BFPS and IPEA, 2025).
Table D.2: 2SLS — Estimates Conditioning on Sex
Homicide Rate
Men Women
(1) (2)
Lagged Enforcement -0.754∗∗ 0.005
(0.343) (0.091)
Average homicide rate across municipalities 25.05 3.02
FE: Municipality & Year
All controls
Observations 5,210 5,210
Municipalities 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) reports results for men and Column (2) for women. “Homicide Rate” is the number of homicide victims of the corresponding sex per 100,000 inhabitants. “Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006-2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
Race
Table D.3 reports the annual distribution of homicide victims by aggregated race groups. The data show that the majority of homicide victims are consistently classified as African Brazilian, representing around 58–69% of all cases over the period, while White and Yellow account for roughly 30–35% and Indigenous individuals for less than 1%. This concentration among the African Brazilian population mirrors the persistent racial disparities in homicide victimization documented by BFPS and IPEA (2025) at the national level.
Table D.3: Annual Totals and Shares of Deaths by Aggregated Race Groups
Levels Shares
Year White and African Indigenous Unknown Total White and African Indigenous Unknown
Yellow Brazilian Yellow (%) Brazilian (%) (%) (%)
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
2006 23,840 38,745 229 4,309 67,123 35.5 57.7 0.3 6.4
2007 23,923 41,231 266 4,319 69,739 34.3 59.1 0.4 6.2
2008 24,757 43,797 270 4,094 72,918 34.0 60.1 0.4 5.6
2009 24,109 44,915 221 4,442 73,687 32.7 61.0 0.3 6.0
2010 23,460 43,560 244 4,129 71,393 32.9 61.0 0.3 5.8
2011 23,074 46,275 299 4,345 73,993 31.2 62.5 0.4 5.9
2012 23,704 48,442 331 4,721 77,198 30.7 62.8 0.4 6.1
2013 23,776 50,737 343 4,667 79,523 29.9 63.8 0.4 5.9
2014 23,244 50,466 379 4,063 78,152 29.7 64.6 0.5 5.2
2015 23,854 53,418 347 3,985 81,604 29.2 65.5 0.4 4.9
2016 24,169 58,727 418 2,073 85,387 28.3 68.8 0.5 2.4
Notes: Columns (2)–(5) report annual totals of deaths for aggregated race groups. Column (6) reports the total number of deaths. Columns (7)–(10) present the share of each aggregated group relative to the total for the corresponding year, expressed in percent. Source: SIM-DataSUS
Table D.4 presents the 2SLS estimates by race group. The results indicate that environmental enforcement significantly reduces homicides among the African Brazilian population: the coefficient equals –0.742 and is statistically significant at the 5% level. Given an average homicide rate of 21.84 per 100,000 inhabitants in this group, the effect implies that the issuance of one additional fine is associated with a reduction of about 3.4% relative to the mean homicide rate. By contrast, the effect is small and statistically insignificant for the White and Yellow population (–0.075) and for Indigenous individuals (0.061). These findings underscore that the deterrent effects of enforcement are concentrated among African Brazilian individuals, the group disproportionately exposed to violence in the Amazon region.
Table D.4: 2SLS — Estimates Conditioning on Race
Homicide Rate
White and Yellow African Brazilian Indigenous
Lagged Enforcement -0.075 -0.742∗∗ 0.061
(0.127) (0.309) (0.060)
Average homicide rate across municipalities 4.49 21.84 1.11
FE: Municipality & Year
All controls
Observations 5,210 5,210 5,210
Municipalities 521 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) reports results for the white and yellow population, Column (2) for the black and mixed-race population, and Column (3) for the indigenous population. ”Homicide Rate” is the number of homicides victims in each municipality per 100,000 inhabitants. ”Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006-2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
Age Groups
Table D.5 reports the annual distribution of homicide victims by age group between 2006 and 2016. The descriptive statistics reveal that the majority of homicide victims are 26 years or older, consistently representing around 59–63% of all cases throughout the period.
Table D.5: Annual Totals and Shares of Deaths by Age Group
Year 0-15 16-25 26-35 36+ Unknown Total Share of Share of Share of Share of Share of
0-15 (%) 16-25 (%) 26-35 (%) 36+ (%) Unknown (%)
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12)
2006 2,226 22,636 17,321 23,412 1,528 67,123 3.3 33.7 25.8 34.9 2.3
2007 2,398 23,409 17,776 24,468 1,688 69,739 3.4 33.6 25.5 35.1 2.4
2008 2,388 23,949 18,991 25,813 1,777 72,918 3.3 32.8 26.0 35.4 2.4
2009 2,419 23,865 19,643 26,122 1,638 73,687 3.3 32.4 26.7 35.4 2.2
2010 2,321 22,824 18,681 26,085 1,482 71,393 3.3 32.0 26.2 36.5 2.1
2011 2,471 23,844 19,480 26,685 1,513 73,993 3.3 32.2 26.3 36.1 2.0
2012 2,515 25,485 19,949 27,700 1,549 77,198 3.3 33.0 25.8 35.9 2.0
2013 2,680 26,602 20,702 28,041 1,498 79,523 3.4 33.5 26.0 35.3 1.9
2014 2,477 25,258 20,120 28,982 1,315 78,152 3.2 32.3 25.7 37.1 1.7
2015 2,515 26,878 20,207 30,779 1,225 81,604 3.1 32.9 24.8 37.7 1.5
2016 2,468 28,767 21,264 31,624 1,264 85,387 2.9 33.7 24.9 37.0 1.5
Notes: Columns (2)–(6) report annual totals of deaths for each age group. Column (7) reports the total number of deaths across these four categories. Columns (8)- (12) present the share of each age group relative to the total for the corresponding year, expressed as a percentage. Source: SIM-DataSUS
Table D.6 presents the 2SLS estimates by age group. The results indicate that environmental enforcement significantly reduces homicides only among individuals whose ages are between 26 and 35 years: the coefficient is –0.35 and statistically significant at the 5% level. Given an average homicide rate of 7.59 per 100,000 inhabitants for this age group, this estimate implies that the issuance of one additional fine is associated with a reduction of about 4.6% relative to the mean homicide rate for this age group. By contrast, the effect for younger individuals is close to zero (0.05 for the 0-15 years group and -0.08 for the 16–25 years group) and not statistically significant.
Table D.6: 2SLS — Estimates Conditioning on Age Group
Homicide Rate
0–15 16–25 26-35 36+
Lagged Enforcement 0.050 -0.083 -0.349∗∗ -0.329
(0.076) (0.143) (0.162) (0.202)
Average homicide rate across municipalities 1.22 8.06 7.59 10.56
FE: Municipality & Year
All controls
Observations 5,210 5,210 5,210 5,210
Municipalities 521 521 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) reports results for the 0–15 age group, Column (2) for ages 16–25, Column (3) for ages 26–35, and Column (4) for ages 36 and above. “Homicide Rate” is the number of homicide victims of the corresponding sex per 100,000 inhabitants. “Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006-2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
Appendix E Controlling for Conservation Policies